The demand curve for a consumer is p-d=b/q d and b are constants. Find the price elastically of demand?
1
Expert's answer
2016-04-11T10:17:04-0400
The demand curve for a consumer is p - d = b/q, where d and b are constants, then p = b/q + d and q = b/(p - d).
The price elastically of demand Ed = p/q*q' = p*(p - d)/b*(-b)/(p - d)^2 = -p/(p - d) = p/(d - p)